# Without Paired Labeled Data: End-to-End Self-Supervised Learning for Drone-view Geo-Localization

### v3

@@ III End-to-End Self-Supervised Learning Method
As illustrated in Fig. 3, the proposed method is composed of two key modules, DHML and ICEL. Additionally, a PLE strategy is employed to optimize the self-supervised learning process. Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with cluster memory and a dual-path contrastive learning strategy to learn intra-view feature representations. To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory. DHML module enables efficient discriminability of intra-view feature representations. Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints. By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form consistent and unified representations.
Problem Formulation: In DVGL, a set of paired drone-satellite images is denoted as $\{x^{d},x^{s}\}$, where $x^{d}$ represents the drone-view image and $x^{s}$ represents the satellite-view image. Unlike the supervised setting, our self-supervised method is developed without using paired drone-satellite images, where the correspondences between drone-view and satellite-view images are entirely unavailable. The objective is to discover reliable cross-view correspondences under such conditions and to accomplish the DVGL task.
@@ III-A Dual-Path Contrastive Learning Baseline
Cluster-based contrastive learning methods have achieved remarkable success in various fields [17, 50]. However, in DVGL, the unique challenges of cross-view scenarios, as discussed in Section I, make it difficult for single-path cluster-based contrastive learning to effectively capture view-specific features. Therefore, we design the baseline network with a dual-path architecture, where drone-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels. Subsequently, a dual-path contrastive learning strategy [51] is employed to enhance intra-view representation learning.
@@
$\begin{split}&\mathbf{f}^{d}_{i}=\mathcal{F}_{\text{backbone}}(x^{d}_{i}),\quad i=1,\dots,N\\ &\mathbf{f}^{s}_{j}=\mathcal{F}_{\text{backbone}}(x^{s}_{j}),\quad j=1,\dots,M\end{split}$ (1)
@@
where $\mathcal{H}_{k}^{d}$ denotes the $k$-th cluster set in drone-view, $k=1,\cdots,K$, and $\mathcal{H}_{l}^{s}$ denotes the $l$-th cluster set in satellite-view, $l=1,\cdots,L$. ||$\cdot$|| ⋅ | indicates the number of instances per cluster.
@@
where $\alpha$ is the momentum updating factor.
@@
$\mathcal{L}^{d}_{c1}=-\log\frac{\exp(q^{d}\cdot\phi^{d}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q^{d}\cdot\phi^{d}_{k}/\tau)}$ (5)
$\mathcal{L}^{s}_{c1}=-\log\frac{\exp(q^{s}\cdot\phi^{s}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q^{s}\cdot\phi^{s}_{l}/\tau)}$ (6)
where $\tau$ denotes the temperature hyper-parameter. Cluster centroids $\phi^{d}_{k}$ and $\phi^{s}_{l}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, $\phi^{d}_{+}$ and $\phi^{s}_{+}$ are positive cluster features associated with $q^{d}$ and $q^{s}$. The final optimization for baseline is denoted as
$\mathcal{L}_{c1}=\mathcal{L}^{d}_{c1}+\mathcal{L}^{s}_{c1}$ (7)
@@ III-B Dynamic Hierarchical Memory Learning
where $\varphi_{1}^{d{{}^{\prime}}},\dots,\varphi_{K}^{d{{}^{\prime}}}$ represent the short-term memory embeddings that encode the cluster representations of the drone-view, and $\varphi_{1}^{d{{}^{\prime\prime}}},\dots,\varphi_{K}^{d{{}^{\prime\prime}}}$ represent the long-term memory embeddings that store historically accumulated stable representations of the drone-view.
Unlike the weight updating scheme in Eq.(4), we assign different weights to the input features and historical memory during the long-term memory update. Specifically, since the momentum parameter is set to 0.2 throughout this study, we enhance the influence of historical memory during the initialization of long-term memory updates. The update procedure can defined as
$\quad\varphi_{k}^{d{{}^{\prime\prime}}}\leftarrow(0.5-\alpha)\varphi_{k}^{d{{}^{\prime\prime}}}+\alpha q^{d{{}^{\prime\prime}}}$ (9)
where $q^{d{{}^{\prime\prime}}}$ represents the drone query instance features. This obtains a new dictionary of long-term memory.
In addition, we further introduce an adaptive weight update mechanism to update the long-term memory. In detail, we compute the euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range $[0,1]$ via the sigmoid function, yielding an adaptive update coefficient:
@@
$\quad\varphi_{k}^{d{{}^{\prime}}}\leftarrow\beta\cdot\varphi_{k}^{d{{}^{\prime\prime}}}+(1-\beta)\cdot\varphi_{k}^{d{{}^{\prime}}}$ (11)
Then, the long-term and short-term memories are fused in a weighted manner to yield the combined feature $\varphi_{k}^{db}$ representation:
@@
Given the drone query $q^{d}$, we compute the contrastive loss by measuring their similarity to the cluster representations maintained in the memory dictionary by the following equations:
@@
$\mathcal{L}_{c3}=\mathcal{L}^{d}_{c3}+\mathcal{L}^{s}_{c3}$ (14)
@@ III-C Information Consistency Evolution Learning
To facilitate the following description, we take the scenario where the drone-view serves as the query as an illustrative example. The similarity between a given drone query $q_{v}^{d}$ that denotes the $v$-th query sample in the drone-view, $v=1,\dots N^{b}$. $N^{b}$ denote the batch size of the query $q_{d}$. Each drone instance $\mathbf{f}_{i}^{d}$ and each satellite instance $\mathbf{f}_{j}^{{s}}$ in the training set are compared to drone query $q_{d}$ by computing a similarity score, respectively. This is defined as
$\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})=\frac{q_{v}^{d}\cdot\mathbf{f}_{i}^{d}}{\|q_{v}^{d}\|_{2}\|\mathbf{f}_{i}^{d}\|_{2}},\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})=\frac{q_{v}^{d}\cdot\mathbf{f}_{j}^{s}}{\|q_{v}^{d}\|_{2}\|\mathbf{f}_{j}^{s}\|_{2}}$ (16)
@@
$\Omega^{dd}=\Big{\{}\mathbf{f}_{i}^{d}\,\Big{|}\,S\big{(}q_{v}^{d},\mathbf{f}_{i}^{d}\big{)}>\gamma\max_{i\in\{1,\dots,N\}}S\big{(}q_{v}^{d},\mathbf{f}_{i}^{d})\Big{\}}$ (17)
$\Omega^{ds}=\Big{\{}\mathbf{f}_{j}^{s}\,\Big{|}\,S\big{(}q_{v}^{d},\mathbf{f}_{j}^{s}\big{)}>\gamma\max_{j\in\{1,\dots,M\}}S\big{(}q_{v}^{d},\mathbf{f}_{j}^{s})\Big{\}}$ (18)
@@
Additionally, we introduce a ranking-based neighborhood selection strategy. Specifically, we independently select the top-$k_{1}$ most relevant samples to construct a strictly filtered neighborhood, ensuring high-confidence associations, while also selecting the top-$k_{2}$ samples define an expanded neighborhood, capturing additional informative relationships. The process can be defined as
$\mathcal{N}_{k_{1}}^{dd}=\Big{\{}\mathbf{f}_{i}^{d}\mid i\in\arg\max_{I\subset\{1,\dots,N\},|I|=k_{1}}\sum_{i\in I}S(q_{v}^{d},\mathbf{f}_{i}^{d})\Big{\}}$ (19)
$\mathcal{N}_{k_{1}}^{ds}=\Big{\{}\mathbf{f}_{j}^{s}\mid j\in\arg\max_{J\subset\{1,\dots,M\},|J|=k_{1}}\sum_{j\in J}S(q_{v}^{d},\mathbf{f}_{j}^{s})\Big{\}}$ (20)
@@
$\mathcal{L}_{\Omega}^{dd}\!=\!-\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{\mathbf{f}_{i}^{d}\in\Omega^{dd}}\!\log\frac{\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{i}^{d})/\tau\right)}{\sum_{\mathbf{f}_{i}^{d}\in{{{\Omega}^{dd}}}}\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{i}^{d})/\tau\right)}$ (21)
$\mathcal{L}_{\Omega}^{ds}\!=\!-\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{\mathbf{f}_{j}^{s}\in\Omega^{ds}}\!\log\frac{\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{j}^{s})/\tau\right)}{\sum_{\mathbf{f}_{j}^{s}\in{{{\Omega}^{ds}}}}\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{j}^{s})/\tau\right)}$ (22)
@@
$p_{v,i}^{dd}=\frac{\exp(\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d}))}{\sum_{\mathbf{f}_{i}^{d}\in\mathcal{N}_{k_{2}}^{dd}}\exp(S(q_{v}^{d},\mathbf{f}_{i}^{d}))}$ (23)
$p_{v,j}^{ds}=\frac{\exp(\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s}))}{\sum_{\mathbf{f}_{j}^{s}\in\mathcal{N}_{k_{2}}^{ds}}\exp(S(q_{v}^{d},\mathbf{f}_{j}^{s}))}$ (24)
$\mathcal{L}_{k_{2}}^{dd}=\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{i\in\mathcal{N}_{k_{2}}^{dd}}p_{v,i}^{dd}\log\left(k_{2}\cdot p_{v,i}^{dd}\right)$ (25)
$\mathcal{L}_{k_{2}}^{ds}=\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{j\in\mathcal{N}_{k_{2}}^{ds}}p_{v,j}^{ds}\log\left(k_{2}\cdot p_{v,j}^{ds}\right)$ (26)
@@
Within a more restrictive and smaller set of $k_{1}$ neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations. This complements the consistency constraint described above and achieves a balance between global structural alignment and local discriminative information. Within the selected $k_{1}$ neighbors, we compute the similarity between the query $q_{v}^{d}$ and each neighbor $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ using the function $S(\cdot,\cdot)$, and apply softmax to obtain the posterior probability for each neighbor. It can be described as
$p^{dd}(i|v)=\frac{\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})\big{)}}{\sum_{\mathbf{f}_{i}^{d}\in\mathcal{N}_{k_{1}}^{dd}}\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})\big{)}}$ (27)
$p^{ds}(j|v)=\frac{\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})\big{)}}{\sum_{\mathbf{f}_{j}^{s}\in\mathcal{N}_{k_{1}}^{ds}}\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})\big{)}}$ (28)
@@
$\mathcal{L}_{k_{1}}^{dd}=-\frac{1}{N_{b}}\sum_{v=1}^{N_{b}}\sum_{i=1}^{k_{1}}p^{dd}(i|v)\log\frac{p^{dd}(i|v)}{1/R}$ (29)
$\mathcal{L}_{k_{1}}^{ds}=-\frac{1}{N_{b}}\sum_{v=1}^{N_{b}}\sum_{j=1}^{k_{1}}p^{ds}(j|v)\log\frac{p^{ds}(j|v)}{1/R}$ (30)
For each view, the loss is formulated as a weighted sum of the original neighborhood loss $\mathcal{L}_{\Omega}$, the strict neighborhood loss $\mathcal{L}_{k_{1}}$, and the extended neighborhood loss $\mathcal{L}_{k_{2}}$. Specifically, the intra-view loss $\mathcal{L}^{dd}$ and the cross-view loss $\mathcal{L}^{ds}$ are respectively defined as
$\mathcal{L}_{c4}^{dd}=\mathcal{L}_{\Omega}^{dd}+\lambda_{1}\mathcal{L}_{k_{1}}^{dd}+\lambda_{2}\mathcal{L}_{k_{2}}^{dd}$ (31)
$\mathcal{L}_{c4}^{ds}=\mathcal{L}_{\Omega}^{ds}+\lambda_{1}\mathcal{L}_{k1}^{ds}+\lambda_{2}\mathcal{L}_{k2}^{ds}$ (32)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the contribution of strict and extended neighborhood components. When the satellate view as the query, the $\mathcal{L}_{c4}^{sd}$ and $\mathcal{L}_{c4}^{ss}$ can be obtained in the same way. The total neighbor learning loss is computed as $\mathcal{L}_{c4}=\mathcal{L}_{c4}^{dd}+\mathcal{L}_{c4}^{ds}+\mathcal{L}_{c4}^{sd}+\mathcal{L}_{c4}^{ss}$.
Finally, the overall loss of our method is defined as
$\mathcal{L}_{total}=\mathcal{L}_{c1}+\mathcal{L}_{c3}+\mathcal{L}_{c4}$ (33)

### v5

@@ III End-to-End Self-Supervised Learning Method
As illustrated in Fig. 2, the proposed method is composed of two key modules, DHML and ICEL. Additionally, a PLE strategy is employed to optimize the self-supervised learning process. Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with the cluster memory and the dual-path contrastive learning strategy to learn intra-view feature representations. To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory. DHML module enables efficient discriminability of intra-view feature representations. Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints. By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form consistent and unified representations.
Problem Formulation: In DVGL, a set of paired drone-satellite images is denoted as $\{x^{d},x^{s}\}$, where $x^{d}$ represents the drone-view image and $x^{s}$ represents the satellite-view image. Unlike the supervised setting, our self-supervised method is developed without using paired drone-satellite images, where the correspondences between drone-view and satellite-view images are entirely unavailable. The objective is to discover reliable cross-view correspondences under such conditions to perform the DVGL task.
@@ III-A Dual-Path Contrastive Learning Baseline
Cluster-based contrastive learning methods have achieved remarkable success in various fields [dai2022cluster, hu2022feature]. However, in DVGL, the unique challenges of cross-view scenarios, as discussed in Section I, make it difficult for single-path cluster-based contrastive learning to effectively capture view-specific features. Therefore, we design the baseline network with a dual-path architecture, where drone-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels. Subsequently, a dual-path contrastive learning strategy [yang2023dual] is employed to enhance intra-view representation learning.
@@
$\begin{split}&\mathbf{f}^{d}_{i}=\mathcal{F}_{\text{backbone}}(x^{d}_{i}),\quad i=1,\dots,N\\ &\mathbf{f}^{s}_{j}=\mathcal{F}_{\text{backbone}}(x^{s}_{j}),\quad j=1,\dots,M.\end{split}$ (1)
@@
where $\mathcal{H}_{k}^{d}$ denotes the $k$-th cluster set in drone-view, $k=1,\cdots,K$, and $\mathcal{H}_{l}^{s}$ denotes the $l$-th cluster set in satellite-view, $l=1,\cdots,L$. $|\cdot|$ indicates the number of instances per cluster.
@@
where $\alpha$ is the momentum factor and is set to 0.1.
@@
$\mathcal{L}^{d}_{c1}=-\log\frac{\exp(q^{d}\cdot\phi^{d}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q^{d}\cdot\phi^{d}_{k}/\tau)}$ (5a)
$\mathcal{L}^{s}_{c1}=-\log\frac{\exp(q^{s}\cdot\phi^{s}_{+}/\tau)}{\sum_{l=0}^{L}\exp(q^{s}\cdot\phi^{s}_{l}/\tau)}$ (5b)
where $\tau$ denotes the temperature hyper-parameter. Cluster centroids $\phi^{d}_{k}$ and $\phi^{s}_{l}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, $\phi^{d}_{+}$ and $\phi^{s}_{+}$ are positive cluster features associated with $q^{d}$ and $q^{s}$. The final optimization for the baseline is denoted as
$\mathcal{L}_{c1}=\mathcal{L}^{d}_{c1}+\mathcal{L}^{s}_{c1}.$ (6)
Remark 1
In our baseline, both views share a common backbone to extract unified feature representations. For each view, features from samples with the high similarity are averaged to create a cluster centroid, which is then used as the representative feature for the cluster. The cluster centroid continuously refined through a momentum-based update strategy as shown in Eq.(4). Building upon this, we minimize the discrepancy between the predicted similarity distribution of query features and the target cluster centroids to guide the query features closer to the target cluster centroids. Simultaneously, we suppress similarities to non-target cluster centroids to enlarge the margin between distinct semantic clusters.
@@ III-B Dynamic Hierarchical Memory Learning
where $\varphi_{1}^{d{{}^{\prime}}},\dots,\varphi_{K}^{d{{}^{\prime}}}$ represent the short-term memory embeddings that encode the cluster representations of the drone-view, and $\varphi_{1}^{d{{}^{\prime\prime}}},\dots,\varphi_{K}^{d{{}^{\prime\prime}}}$ represent the long-term memory embeddings that store historically accumulated stable representations of the drone-view. Since the clustering representations in the short-term and long-term memories adopt a different value of momentum factor from that in Eq.(2) due to different motivations, their update dynamics and resulting representations are also different. Therefore, we use different symbols to distinguish them.
Unlike Eq.(4), we assign different momentum factor to represent the weights between the input features and historical memory during the long-term memory update. Specifically, since the momentum parameter $\xi$ is set to 0.3, the influence of historical memory during the long-term memory updates has been enhanced. The updating procedure can be defined as
$\quad\varphi_{k}^{d{{}^{\prime\prime}}}\leftarrow(1-\xi)\varphi_{k}^{d{{}^{\prime\prime}}}+\xi q^{d{{}^{\prime\prime}}}$ (8)
where $q^{d{{}^{\prime\prime}}}$ represents the drone query instance feature in long-term memory. This obtains a new dictionary of long-term memory.
In addition, we also introduce an adaptive weight update mechanism to update the short-term memory. In detail, we compute the euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range $[0,1]$ via the sigmoid function, which yields an adaptive update coefficient
@@
$\quad\varphi_{k}^{d{{}^{\prime}}}\leftarrow\beta\cdot\varphi_{k}^{d{{}^{\prime\prime}}}+(1-\beta)\cdot\varphi_{k}^{d{{}^{\prime}}}.$ (10)
Then, the long-term and short-term memories are fused in a weighted manner to yield the combined feature $\varphi_{k}^{db}$ representation,
@@
Given the drone query $q^{d}$, we compute the contrastive loss by measuring their similarity to the cluster representations maintained in the memory dictionary by the following equations,
@@
$\mathcal{L}_{c3}=\mathcal{L}^{d}_{c3}+\mathcal{L}^{s}_{c3}.$ (13)
Remark 2
The proposed DHML module differs from traditional single-memory methods [dai2022cluster, adca, yang2024shallow]. DHML is modeled as a dual-cycle process across two temporal scales. The long-term memory accumulates stable global features over a long period and is updated with a small momentum to maintain consistency, while the short-term memory rapidly responds to variations in the current input over a short period to capture local dynamic features. The adaptive coefficient $\beta$ functions as a temporal regulator between the two memories. When the input features deviate significantly from the long-term memory, $\beta$ increases and the short-term memory relies more on the stable long-term representation. When the deviation is small, $\beta$ decreases, allowing the short-term memory to retain more of its recent, immediate information.
@@ III-C Information Consistency Evolution Learning
To facilitate the following description, we take the scenario where the drone-view serves as the query as an illustrative example. The similarity between a given drone query $q_{v}^{d}$ that denotes the $v$-th query sample in the drone-view, $v=1,\dots N^{b}$. $N^{b}$ denote the batch size of the query $q_{d}$. Each drone instance $\mathbf{f}_{i}^{d}$ and each satellite instance $\mathbf{f}_{j}^{{s}}$ in the training set are compared with drone query $q_{d}$ by computing a similarity score, respectively. This is defined as
$\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})=\frac{q_{v}^{d}\cdot\mathbf{f}_{i}^{d}}{\|q_{v}^{d}\|_{2}\|\mathbf{f}_{i}^{d}\|_{2}},\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})=\frac{q_{v}^{d}\cdot\mathbf{f}_{j}^{s}}{\|q_{v}^{d}\|_{2}\|\mathbf{f}_{j}^{s}\|_{2}}.$ (15)
@@
$\Omega^{dd}=\Big\{\mathbf{f}_{i}^{d}\,\Big|\,S\big(q_{v}^{d},\mathbf{f}_{i}^{d}\big)>\gamma\max_{i\in\{1,\dots,N\}}S\big(q_{v}^{d},\mathbf{f}_{i}^{d})\Big\}$ (16a)
$\Omega^{ds}=\Big\{\mathbf{f}_{j}^{s}\,\Big|\,S\big(q_{v}^{d},\mathbf{f}_{j}^{s}\big)>\gamma\max_{j\in\{1,\dots,M\}}S\big(q_{v}^{d},\mathbf{f}_{j}^{s})\Big\}$ (16b)
@@
Additionally, we introduce a ranking-based neighborhood selection strategy. Specifically, we independently select the top-$k_{1}$ most relevant samples to construct a strictly filtered neighborhood, ensuring high-confidence associations, while also selecting the top-$k_{2}$ samples define an expanded neighborhood, capturing additional informative relationships. Note that $k_{2}>k_{1}$. Here $\mathrm{TopK}_{S}(q,\mathcal{F},k)$ denotes the ranking-based neighborhood selection strategy, which returns the top-$k$ elements in $\mathcal{F}$ ranked by their similarity to $q$. The process can be defined as
$\mathcal{N}_{k_{1}}^{dd}$ $=\mathrm{TopK}_{S}\!\big(q_{v}^{d},\{\mathbf{f}_{i}^{d}\}_{i=1}^{N},k_{1}\big),$ (17a)
$\mathcal{N}_{k_{1}}^{ds}$ $=\mathrm{TopK}_{S}\!\big(q_{v}^{d},\{\mathbf{f}_{j}^{s}\}_{j=1}^{M},k_{1}\big),$ (17b)
@@
$\mathcal{L}_{\Omega}^{dd}\!=\!-\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{\mathbf{f}_{i}^{d}\in\Omega^{dd}}\!\log\frac{\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{i}^{d})/\tau\right)}{\sum_{\mathbf{f}_{i}^{d}\in{{{\Omega}^{dd}}}}\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{i}^{d})/\tau\right)}$ (18a)
$\mathcal{L}_{\Omega}^{ds}\!=\!-\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{\mathbf{f}_{j}^{s}\in\Omega^{ds}}\!\log\frac{\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{j}^{s})/\tau\right)}{\sum_{\mathbf{f}_{j}^{s}\in{{{\Omega}^{ds}}}}\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{j}^{s})/\tau\right)}$ (18b)
@@
$p_{v,i}^{dd}=\frac{\exp(\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d}))}{\sum_{\mathbf{f}_{i}^{d}\in\mathcal{N}_{k_{2}}^{dd}}\exp(S(q_{v}^{d},\mathbf{f}_{i}^{d}))}$ (19a)
$p_{v,j}^{ds}=\frac{\exp(\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s}))}{\sum_{\mathbf{f}_{j}^{s}\in\mathcal{N}_{k_{2}}^{ds}}\exp(S(q_{v}^{d},\mathbf{f}_{j}^{s}))}$ (19b)
$\mathcal{L}_{k_{2}}^{dd}=\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{i\in\mathcal{N}_{k_{2}}^{dd}}p_{v,i}^{dd}\log\left(k_{2}\cdot p_{v,i}^{dd}\right)$ (19c)
$\mathcal{L}_{k_{2}}^{ds}=\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{j\in\mathcal{N}_{k_{2}}^{ds}}p_{v,j}^{ds}\log\left(k_{2}\cdot p_{v,j}^{ds}\right)$ (19d)
@@
Within a more restrictive and smaller set of $k_{1}$ neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations. This complements the consistency constraint described above and achieves a balance between global structural alignment and local discriminative information. Within the selected $k_{1}$ neighbors, we compute the similarity between the query $q_{v}^{d}$ and each neighbor $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ using the function $\mathcal{S}(\cdot,\cdot)$, and apply softmax to obtain the posterior probability for each neighbor. It can be described as
$p^{dd}(i|v)=\frac{\exp\big(\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})\big)}{\sum_{\mathbf{f}_{i}^{d}\in\mathcal{N}_{k_{1}}^{dd}}\exp\big(\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})\big)}$ (20a)
$p^{ds}(j|v)=\frac{\exp\big(\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})\big)}{\sum_{\mathbf{f}_{j}^{s}\in\mathcal{N}_{k_{1}}^{ds}}\exp\big(\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})\big)}.$ (20b)
@@
$\mathcal{L}_{k_{1}}^{dd}=-\frac{1}{N_{b}}\sum_{v=1}^{N_{b}}\sum_{i=1}^{k_{1}}p^{dd}(i|v)\log\frac{p^{dd}(i|v)}{1/R}$ (21a)
$\mathcal{L}_{k_{1}}^{ds}=-\frac{1}{N_{b}}\sum_{v=1}^{N_{b}}\sum_{j=1}^{k_{1}}p^{ds}(j|v)\log\frac{p^{ds}(j|v)}{1/R}.$ (21b)
For each view, the loss is formulated as a weighted sum of the original neighborhood loss $\mathcal{L}_{\Omega}$, the strict neighborhood loss $\mathcal{L}_{k_{1}}$, and the extended neighborhood loss $\mathcal{L}_{k_{2}}$. Specifically, the intra-view loss $\mathcal{L}^{dd}$ and the cross-view loss $\mathcal{L}^{ds}$ are, respectively, defined as
$\mathcal{L}_{c4}^{dd}=\mathcal{L}_{\Omega}^{dd}+\lambda_{1}\mathcal{L}_{k_{1}}^{dd}+\lambda_{2}\mathcal{L}_{k_{2}}^{dd}$ (22a)
$\mathcal{L}_{c4}^{ds}=\mathcal{L}_{\Omega}^{ds}+\lambda_{1}\mathcal{L}_{k_{1}}^{ds}+\lambda_{2}\mathcal{L}_{k_{2}}^{ds}$ (22b)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the contribution of strict and extended neighborhood components. When using the satellite-view as the query, $\mathcal{L}_{c4}^{sd}$ and $\mathcal{L}_{c4}^{ss}$ can be obtained in the same way. The total neighbor learning loss is computed as $\mathcal{L}_{c4}=\mathcal{L}_{c4}^{dd}+\mathcal{L}_{c4}^{ds}+\mathcal{L}_{c4}^{sd}+\mathcal{L}_{c4}^{ss}$.
Finally, the overall loss of DMNIL is defined as
$\mathcal{L}_{total}=\mathcal{L}_{c1}+\mathcal{L}_{c3}+\mathcal{L}_{c4}.$ (23)
Remark 3
The ICEL module computes the association distribution between query samples and all candidate samples using a similarity measure $\mathcal{S}(\cdot,\cdot)$ and constructs multi-scale neighborhood structures through threshold-based and ranking-based selection. For a drone query, this includes a high-confidence distribution neighborhood $\Omega^{dd}$ and $\Omega^{ds}$, an extended neighborhood $\mathcal{N}_{k_{2}}^{dd}$ and $\mathcal{N}_{k_{2}}^{ds}$, and a strict neighborhood $\mathcal{N}_{k_{1}}^{dd}$ and $\mathcal{N}_{k_{1}}^{ds}$. Within the extended neighborhood, a normalized similarity distribution is computed and optimized by minimizing entropy. In the strict neighborhood, a conditional probability distribution between the query and its neighbors is established and the KL divergence to a uniform distribution is minimized. Satellite queries are processed in the same manner.